A manager bought 12 pounds of peanuts for $30. He wants to mix $5 per pound cashews with the peanuts to get a batch of mixed nuts that is worth $4 per pound. How many pounds of cashews are needed?
Answer by checkley77(12844)

5X+12*2.5=4(12+X)
5X+30=48+4X
5X-4X=48-30
X=18 POUNDS OF CASHEWS IS NEEDED.
PROOF:
5*18+30=4(12+18)
90+30=4*30
120=120